# Bootstrapping Classical Shadows for Neural Quantum State Tomography

### v2

@@ III.3 Applications on Mixed GHZ States
In this section, we describe our application of the methods introduced earlier, namely, the shadow-based cross-entropy loss function with stabilizer-based sampling, to mixed quantum states generated by noisy circuits. Specifically, we consider a circuit that prepares a six-qubit GHZ state beginning from the all-zeroes state |$\ket{0}^{\otimes 6}$. The circuit is composed of a Hadamard gate applied to the first qubit, followed by a series of CNOT gates applied to the consecutive pairs of qubits. For the noise model, we assume two-qubit depolarization noise on the CNOT gates. The depolarizing noise channel $\rho\mapsto(1-\frac{16}{15}p)\rho+\frac{16}{15}p\mathbf{I}/4$ is parameterized via a noise strength parameter $p$. Motivated by the experimental relevance of our technique, we consider Pauli measurement datasets, as those are less noisy than Clifford measurements. As the ansatz for the density matrix, we employ the purification ansatz Eq. (2) with a transformer neural network used to parameterize $\psi_{\lambda}(s,\bar{s})$. This approach allows us to impose physical constraints on the ansatz.
@@ Appendix A The Transformer Neural Quantum State and Details of Numerical Experiments
We base our neural quantum states on the transformer architecture from Ref. [3] (for details on its implementation, refer to Section I of the supplementary material of that work). More specifically, a quantum state is represented using the transformer model parameterized by the neural network weights $\lambda$ as specified in Eq. 1 or Eq. 2, depending on whether the target quantum state is a pure state or a mixed state, respectively. The details of the transformer’s architecture, hyperparameters, and datasets are presented in Table 1. PyTorch [15] is used for the implementation and training of the transformer. The Stim stabilizer-circuit simulator [16] is used for dataset generation and classical shadow state reconstruction. For training, we use the Adam optimizer [17] for stochastic-gradient descent. A cosine annealing schedule is applied to systematically adjust the learning rate throughout training epochs [18]. Finally, early stopping is used to prevent over-fitting.
@@
Table 1: Numerical experiment settings for Figs. 3, 4 and 5.

### v3

@@ III.3 Applications on Mixed GHZ States
In this section, we describe our application of the methods introduced earlier, namely, the shadow-based cross-entropy loss function with stabilizer-based sampling, to mixed quantum states generated by noisy circuits. Specifically, we consider a circuit that prepares a six-qubit GHZ state beginning from the all-zeroes state |$\ket{0}^{\otimes 6}$. The circuit is composed of a Hadamard gate applied to the first qubit, followed by a series of CNOT gates applied to the consecutive pairs of qubits. For the noise model, we assume two-qubit depolarization noise on the CNOT gates. The depolarizing noise channel $\rho\mapsto(1-\frac{16}{15}p)\rho+\frac{16}{15}p\mathbb{I}/4$ is parameterized via a noise strength parameter $p$. Motivated by the experimental relevance of our technique, we consider Pauli measurement datasets, as those are less noisy than Clifford measurements. As the ansatz for the density matrix, we employ the purification ansatz Eq. (2) with a transformer neural network used to parameterize $\psi_{\lambda}(s,\bar{s})$. This approach allows us to impose physical constraints on the ansatz.
@@ Appendix A The Transformer Neural Quantum State and Details of Numerical Experiments
We base our neural quantum states on the transformer architecture from Ref. [bennewitz2021] (for details on its implementation, refer to Section I of the supplementary material of that work). More specifically, a quantum state is represented using the transformer model parameterized by the neural network weights $\lambda$ as specified in Eq. 1 or Eq. 2, depending on whether the target quantum state is a pure state or a mixed state, respectively. The details of the transformer’s architecture, hyperparameters, and datasets are presented in Table 1. PyTorch [16] is used for the implementation and training of the transformer. The Stim stabilizer-circuit simulator [17] is used for dataset generation and classical shadow state reconstruction. For training, we use the Adam optimizer [18] for stochastic-gradient descent. A cosine annealing schedule is applied to systematically adjust the learning rate throughout training epochs [19]. Finally, early stopping is used to prevent over-fitting.
@@
Table 1: Numerical experiment settings for Figs. 3, 4, and 5.